OpenStudy (anonymous):

***WILL MEDAL*** Find the length of an arc of 40° in a circle with an 8 inch radius. a) 8pi/9 inches b) 16pi/9 inches c) 64pi/9 inches

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

ok so first we need the arc length formula which is \[s = \theta r \] where s = arc length, \[\theta \] = angle , and r is the radius.

OpenStudy (anonymous):

Oh wow they never gave me a formula for this

OpenStudy (usukidoll):

so we are given radius = 8 and our angle is 40 degrees, but I need that in radians so I need to convert \[40 \times \frac{\pi}{180}\] \[\frac{40 \pi}{180}\] \[\frac{2\pi}{9}\]

OpenStudy (usukidoll):

I think we almost got our arc length \[s= \frac{2\pi}{9} \times 8\]

OpenStudy (usukidoll):

just gotta multiply what's 2pi x 8 ?

OpenStudy (anonymous):

Wait a sec I need to get my calculator

OpenStudy (anonymous):

brb

OpenStudy (usukidoll):

actually... we can just leave the pi part alone ._. and just calculate 2 x 8 for the numerator.

OpenStudy (anonymous):

Sorry. I'm back now

OpenStudy (anonymous):

2x8=16

OpenStudy (usukidoll):

yes so our arc length is \[s= \frac{16 \pi}{9}\]

OpenStudy (anonymous):

So that makes the answer B

OpenStudy (anonymous):

Thank you so much. Now I'm sure I know how to do this. Especially since now I know what the formula is.

OpenStudy (usukidoll):

hmmm... Why weren't you given the formula? It was supposed to be given to you x_X

OpenStudy (anonymous):

I know, but they didn't. At least, it didn't show it on the lesson.

OpenStudy (anonymous):

I'll be right back. It'll be a few minutes

OpenStudy (anonymous):

gtg

OpenStudy (usukidoll):

k bye... wow shame on whoever for not giving it. -_-!

OpenStudy (anonymous):

I'm back